OpenStudy (dancergirl45):

Which exponential function goes through the points (1, 16) and (4, 128)? f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)−x f(x) = 4(4)−x

OpenStudy (dancergirl45):

@DanJS

OpenStudy (dancergirl45):

@AloneS

OpenStudy (dancergirl45):

I think b

alones (alones):

so you think it's c?/

alones (alones):

ooops i mean b xD

alones (alones):

eah I believe your right !!! :)

OpenStudy (dancergirl45):

yes

OpenStudy (dancergirl45):

Thank you

alones (alones):

xD YOU WELCOME :)

OpenStudy (danjs):

Tthis one?

OpenStudy (danjs):

since they actually gave you choices, you can just test which one works for those points, but, there is a way you probably learned to solve for the constants in y = A*B^X 2 data points given for (x,y) can set up 2 equations in terms of A and B

alones (alones):

so is it right?? @DanJS or maybe I made a mistake??

OpenStudy (danjs):

y = a * b^x when x=0, y= a each increase in x by 1 will multiply 'a' by the 'b' base.

OpenStudy (danjs):

x= 0 , 1 .... y = a, a*b , a*b*b, a*b*b*b... exe